# Hal-Eval: A Universal and Fine-grained Hallucination Evaluation Framework for Large Vision Language Models

### v1

@@ 4.3 Utilizing Hal-Data for Supervised Fine-tuning
To validate whether Hal-Data can assist LVLM in eliminating hallucinations and improving performance through instructional fine-tuning, we conducted the following experiment: we constructed instructional data from the high-quality data labeled for GPT-4, totaling 130K, and after combining this instructional data with that of LLaVA1.5, we conducted joint fine-tuning training and get the varient of LLaVA1.5 named Hal-VL. As shown in Table 4 , we evaluate Hal-VL on Hal-Eval and achieve the best performance which demonstrate that Hal-Data can aid in the elimination of hallucinations within LVLMs. In addition to evaluations on Hal-Eval, we also other general benchmarks such as VQA Goyal et al. (2017), GQA Hudson and Manning (2019), MMBench Liu et al. (2023f), MME Fu et al. (2023). Seed Bench Li et al. (2023b) and so on. As shown in Figure 7, the results demonstrate that Hal-Data can also enhances their performance across general benchmarks. (Refer to Appendix C.1 for more results.)
5 Related Work
Large Vision Language Model: Unlike earlier vision-language pre-training models Li et al. (2021, 2022); Jiang et al. (2023e, b, 2022, c, d), that are relatively small in scale, based on LLMs, there are three principal approaches to constructing LVLMs, all demonstrating potential for robust zero-shot generalization in the vision-language field . Flamingo Alayrac et al. (2022) leads the way by combining a static vision encoder with a large language model that has special cross-attention features for linking images and text. On the other hand, PaLM-E Driess et al. (2023) merges visual features into the already powerful PaLM Chowdhery et al. (2022) and this method is also used by models such as LLaVA Liu et al. (2023d) and Shikra Chen et al. (2023a). However, creating long visual sequences is a limitation here. BLIP-2 Li et al. (2023c) addressed this by shortening visual feature sequences with a method inspired by DETR Carion et al. (2020), and this improvement is reflected in subsequent models like Kosmos-1 Huang et al. (2023), mPLUG-Owl Ye et al. (2023).
Hallucination Evaluation: LVLM hallucination benchmarks specifically aim at non-hallucinatory generation or hallucination discernment. These benchmarks are classified according to the type of evaluation approach they follow: Discriminative (Dis) or Generative (Gen). POPE Li et al. (2023e), NOPE Lovenia et al. (2023), and CIEM Hu et al. (2023) are examples of discriminative benchmarks. Each of these benchmarks exclusively directs attention towards object hallucinations and utilizes accuracy as their primary evaluation metric. The metric is calculated by querying the presence of objects within images and comparing the model’s responses with the ground truth. More recent research predominantly accentuates generative benchmarks over discriminative ones. Whilst discriminative benchmarks focus mainly on object-level hallucinations, generative benchmarks widen their scope to encompass a more extensive range of hallucinations, such as attribute and relation hallucinations Gunjal et al. (2023); Liu et al. (2023a); Jing et al. (2023); Wang et al. (2023b); Sun et al. (2023b); Jiang et al. (2023a).
6 Conclusion
@@
Limitation
We acknowledge several limitations within our study. Firstly, the annotations for hallucination types, while extensive, may not encapsulate every possible variant, suggesting room for expansion in terms of the dataset’s coverage. Secondly, the Hal-Eval framework, although comprehensive, may still benefit from further refinement to encompass emerging hallucination patterns as LVLMs continue to evolve. Lastly, while efforts were made to reduce biases within LVLMs using our dataset, the potential for residual biases remains, necessitating continuous monitoring and updating of these models. Future work should focus on addressing these limitations by expanding the scope of hallucination types, refining evaluation frameworks, and further exploring the nuances of hallucination content for the continued betterment of LVLMs.
Ethics Statement
The datasets used for constructing Hal-Data include COCO Lin et al. (2014), Conceptual Captions (CC) Changpinyo et al. (2021), SBU Ordonez et al. (2011), LAION Schuhmann et al. (2022), and ShareGPT4v Chen et al. (2023b). These are all publicly available datasets, and we have not utilized any private datasets in our research. We have obtained the unanimous consent and permission from all annotators involved in the human relevance evaluation experiments. Our methodology is not specifically engineered to filter out potentially harmful or biased image descriptions; furthermore, these biases may also be prevalent in the models that constitute part of our pipeline.
Appendix A Dataset Analysis for Hal-Data
A.1 Data source
@@
For the compilation of the Hal-Data 130K dataset, data were sourced from several established datasets, including: (1) COCO, (2) Conceptual Captions (CC), (3) SBU, (4) LAION, and (5) Share-GPT4. Table 5 and Table 6 provides the detailed distribution ratio of these sources. In the scale-up process of Hal-Data to 2 million entries, priority was given to incorporating 558,000 data points from the LLaVA dataset, which constitutes a curated collection derived from LAION, CC, and SBU. The remaining 1400K entries were likewise obtained from this trio of datasets, ensuring a consistent and diverse array of data for robust model training. As shown in Table 6, we also compared with other hallucinatory datasets like M-HalDetectGunjal et al. (2023) and HaELM Wang et al. (2023a).
A.2 Visualization
as shown in Figure 11, we tokenized the annotations of hallucination content in Hal-Data 130K, and visualized them via a word cloud. The results shown in the word cloud align with the definitions of the hallucination content. For instance, the words in the word cloud for object hallucinations are distinguishably nouns, and for attribute hallucinations, they largely consist of features such as colors. This supports the reliability and validity of our labeled dataset.
Appendix B Experiments Settings
B.1 Training Setting for Hal-Evaluator and Hal-VL
@@
We followed the original approach of LLaVA 1.5 Liu et al. (2023b), we used the complete pre-training dataset of LLaVA 1.5 during the first stage of pre-training. We also keeping the same hyperparameter settings with LLaVA 1.5 . Our experiments were conducted using 16 NVIDIA A100 GPUs with 80G of memory. We used Deepspeed Rajbhandari et al. (2020) for Hal-Evaluator and Hal-VL, with a batch size of 64 on a single GPU.
B.2 Setting for GPT-4 Annotation of AFHA
@@
B.3 Setting for GPT-4 Filtering of AFHA
@@
B.4 Experiment Setting of Chain-of-Thought for Discriminative Evaluation.
@@
Table 7: Evaluation results for different MLLMs on MMHal-Bench.
Dataset Type Model Accuracy Precision Recall F1 Yes (%)
In-domain Object LLaVA 1.5 62.2 76.1 35.6 48.5 23.3
Hal-VL 72.1 89.1 58.3 70.3 40.4
Attribute LLaVA 1.5 62.2 76.1 35.6 48.5 23.3
Hal-VL 72.1 83.2 60.3 69.9 38.2
Relation LLaVA 1.5 55.4 59.1 35.6 44.4 22.1
Hal-VL 66.0 79.1 60.2 68.3 28.2
Event LLaVA 1.5 62.7 77.9 35.6 48.9 22.8
Hal-VL 74.2 86.2 55.3 70.22 26.1
Out-of-domain Object LLaVA 1.5 59.2 86.2 21.9 35.0 18.2
Hal-VL 78.0 94.4 59.6 73.1 37.6
Attribute LLaVA 1.5 58.1 79.4 21.9 34.4 13.8
Hal-VL 77.9 93.9 59.6 72.9 31.7
Relation LLaVA 1.5 53.7 60.2 21.9 32.2 12.7
Hal-VL 72.0 79.2 59.6 68.0 31.6
Event LLaVA 1.5 57.7 77.2 21.9 34.2 14.2
Hal-VL 78.4 95.5 59.6 73.4 31.2
Table 8: Results of LLaVA 1.5 and Hal-VL models under discriminative evaluation of four hallucination types on the in-domain dataset and out-of-domain dataset. Yes denotes the proportion of answering “Yes” to the given question.
@@
Table 9: Generative Hallucination Evaluation for Hal-VL and LLaVA 1.5.
Appendix C More Experiments
C.1 More result of Hal-VL.
In this section, we provide detailed evaluations of Hal-VL on more general benchmarks and Hal-Eval (Table 8 and Table 9). As indicated in the Table 10 and Table 11, we found that Hal-VL achieved significant advantages on the vast majority of these benchmarks. This demonstrates its robustness and generalizability across different tasks and datasets, underlining its effectiveness in handling hallucination problems in LVLMs.
General VQA General VQA (Zero-shot)
Method #Params VQAv2 GQA VizWizQA TextVQA SciQA (IMG)
BLIP-2 Li et al. (2023d) 8.2B 65.0 41.0 19.6 42.5 61.0
InstructBLIP Dai et al. (2023) 8.2B - 49.2 34.5 $50.1^{\dagger}$ 60.5
@@ Appendix C More Experiments
PaLM-E-12B 12B 76.2 - - - -
Shikra 7.2B 77.4 - - - -
Qwen-VL-Chat Bai et al. (2023b) 9.6B 78.2 57.5 38.9 $61.5^{\ddagger}$ 68.2
LLaVA Liu et al. (2023e) 7.2B 71.3 41.3 36.7 $50.2^{\dagger}$ 61.5
MiniGPT-4 Zhu et al. (2023) 7.2B 65.2 30.8 30.2 $52.3^{\dagger}$ 58.4
LLaVA1.5 Liu et al. (2023c) 7.2B 78.5 62.0 50.0 $58.2^{\dagger}$ 66.8
[…]
Table 10: Performance comparison on visual question answering. For VQA, accuracy is reported. Note that specialists are fine-tuned on each individual dataset. † denotes OCR inputs are utilized. ‡ indicates the model has trained on the dataset. We gray out those specialists’ methods which are individually fine-tuned on the dataset as well as those fine-tuned results of generalists.
Figure 8: Evaluation results on Hal-Eval-Dis after fine-tuning Hal-Evaluator using SFT data of varying sizes. We have calculated the average F1 value across all types of hallucinations.
C.2 Impact of SFT Data Scales for Hal-Evaluator
Method MME MMBench MM-Vet SEED-Bench
BLIP-2 Li et al. (2023d) 1293.84 - 22.4 46.4
mPLUG-Owl Ye et al. (2023) 967.34 46.6 - 34.0
InstructBLIP Dai et al. (2023) 1212.82 36.0 26.2 53.4
Otter Li et al. (2023a) 1292.26 48.3 24.6 32.9
Qwen-VL-Chat Bai et al. (2023a) 1487.58 60.6 - 58.2
LLaVA Liu et al. (2023e) 502.82 36.2 28.1 33.5
MiniGPT-4 Zhu et al. (2023) 581.67 23.0 22.1 42.8
LLaVA-1.5 Liu et al. (2023c) 1510.70 64.3 30.5 58.6
Hal-VL 1620.10 66.3 31.3 59.4
Table 11: Zero-shot multi-modal evaluation on multi-modal benchmarks including MME Fu et al. (2023), MMBench Liu et al. (2023f), MM-Vet Yu et al. (2023), SEED-Bench Li et al. (2023b). The overall scores are reported for evaluation. For MMBench, we report test results.
@@ C.2 Impact of SFT Data Scales for Hal-Evaluator
POPE Li et al. (2023e) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
NOPE Lovenia et al. (2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
CIEM Hu et al. (2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
M-HalDetect Gunjal et al. (2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
GAVIE Liu et al. (2023a) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
FAITHScore Jing et al. (2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
HaELM Wang et al. (2023b) $\times$ $\times$ $\times$ $\times$ - - - - $\times$
MMHal-Bench Sun et al. (2023b) $\times$ $\times$ $\times$ $\times$ - - - - $\times$
AMBER Wang et al. (2023a) $\times$ $\times$ $\times$ $\times$
@@
Table 12: Comparison of Hallucination Evaluation Benchmarks for LVLMs.
C.3 Effectiveness of Hall-evaluator for Mitigating Generative Hallucination.
The evaluation model Hal-Evaluator, utilized for generative hallucination assessment, can not only detect and evaluate model hallucinations but also modify the hallucinatory content within model outputs to aid in hallucination elimination. To validate the effectiveness of Hal-Evaluator in eliminating hallucinations, we conducted the following experiment: we used the output from MiniGPT-4 and LLaVA as the input for Hal-Evaluator, allowing Hal-Evaluator to detect and rectify any hallucinatory content. Subsequently, we re-evaluated the corrected output with MMHal-Bench Sun et al. (2023b) as shown in Table 7. Our findings indicate that Hal-Evaluator can effectively eliminate model hallucinations.
C.4 Compared with Other Hallucination Benchmarks
@@
Figure 9: The image shows the prompt we used to elicit hallucination annotations from GPT-4.
Figure 10: The figure illustrates the tailored prompt designed to guide GPT-4 in distinguishing and discarding labels that are inconsistent with the predefined criteria of hallucination types or are marked inaccurately.
Figure 11: We tokenized the annotations of hallucination content in Hal-Data 130K, and visualized them via a word cloud.
C.5 Demo for Hal-Evaluator

### v2

@@ 5.4. Utilizing Hal-Data for Supervised Fine-tuning
To validate whether Hal-Data can assist LVLM in eliminating hallucinations and improving performance through instructional fine-tuning, we conducted the following experiment: we constructed instructional data from Hal-Data 130K, and after combining this instructional data with that of LLaVA1.5, we conducted joint fine-tuning training and get the varient of LLaVA1.5 named Hal-VL.
Hallucination Benchmark: As shown in Table 5 , we evaluate Hal-VL on Hal-Eval and achieve the best performance. In addition to evaluations on Hal-Eval, as shown in Table 9, we also evaluate Hal-VL with other hallucination benchmark like MMHal-Bench (Stiennon et al., 2020). The experiments demonstrate that Hal-VL significantly outperforms LLaVA 1.5. These results indicate Hal-Data effectively aid models in mitigating hallucinations.
General benchmarks: We also evaluate Hal-VL on multiple general benchmark such as VQA (Goyal et al., 2017), GQA (Hudson and Manning, 2019) and so on. As indicated in the Table 6, we found that Hal-VL achieved significant advantages on the vast majority of these benchmarks which suggesting Hal-Data can not only mitigating hallucinations but also enhances the overall model performance. These results support that:
Finding 4: The hallucinatory samples used to train our evaluator also serve as effective supervised fine-tuning data for LVLMs, contributing to the reduction of hallucinations and enhancement of their benchmark performance.
6. Conclusion
@@
Appendix A Automatic Fine-grained Hallucination Annotation Pipline (AFHA)
The existing multimodal hallucination research suffers from a lack of large-scale datasets with fine-grained annotations specific to hallucinations. To address this issue, we design an automatic fine-grained hallucination annotation pipline featuring annotations for four hallucination types and specific hallucination content. Data Annotation. We annotated image-text paired data using the GPT-4 prompt method. We initially established a rigorous definition for various types of hallucinations. Building upon this groundwork, we engaged GPT-4 to rephrase the collated image-text pairs in line with the diverse classifications of hallucinations. This step involved inject distinctive hallucinatory elements into the original captions. An example of a prompt designed to generate ’object’ annotations is illustrated in Figure 7(b). The outcome of this procedure was a collection of image descriptions enriched with specified hallucination categories. Moreover, we delegated to GPT-4 the responsibility of concocting specific hallucinatory content. Consequently, this strategy yielded an extensively annotated dataset facilitated by GPT-4, with samples of annotations across different types of hallucinations displayed in Figure 7(c). For more details, please refer to the Appendix C.2. Data Filtering. Following the initial annotation phase, we identified that the quality of the labeled data remained unsatisfactory. Random sampling revealed that approximately 30% of the annotated dataset still harbored noise that failed to meet our stringent labeling criteria. Hence, we proceeded to craft a tailored prompt to commission GPT-4 for the task of purging and refining the noisy annotations, a process thoroughly outlined in the Appendix C.3. Subsequent to GPT-4’s meticulous cleanup operation, a manual verification process ascertained that over 97% of the data accorded with the stipulated annotation standards.
Appendix B Dataset Analysis for Hal-Data
Name COCO LCS ShareGPT4V
Hal-Data 130K 40% 40% 20%
Hal-Data 2M 0% 100 % 0%
Table 7. The proportion of data sources in Hal-Data 130K and Hal-Data 2M.
Name Data Source Visible Annotated by Samples
M-HalDetect COCO Human 4K
HaELM COCO Human,GPT3.5 15K
Hal-Data 130K LCS, COCO, ShareGPT4V GPT4 130K
Hal-Data 2M LCS Hal-Annotator 1958K
Table 8. Comparison of hallucinatory datasets and Hal-Data. ‘LCS’ abbreviates the LAION (Schuhmann et al., 2022), CC (Changpinyo et al., 2021), and SBU (Ordonez et al., 2011) datasets. The ‘Visible’ column denotes the image visibility during captioning, and the last column shows the average character number of the caption.
B.1. Impact of SFT Data Scales for Hal-Evaluator
B.2. Data source
@@
For the compilation of the Hal-Data 130K dataset, data were sourced from several established datasets, including: (1) COCO, (2) Conceptual Captions (CC), (3) SBU, (4) LAION, and (5) Share-GPT4. Table 7 and Table 8 provides the detailed distribution ratio of these sources. In the scale-up process of Hal-Data to 2 million entries, priority was given to incorporating 558,000 data points from the LLaVA dataset, which constitutes a curated collection derived from LAION, CC, and SBU. The remaining 1400K entries were likewise obtained from this trio of datasets, ensuring a consistent and diverse array of data for robust model training. As shown in Table 8, we also compared with other hallucinatory datasets like M-HalDetect(Gunjal et al., 2023) and HaELM (Wang et al., 2023a).
Appendix C Experiments Settings
C.1. Training Setting for Hal-Evaluator and Hal-VL
@@
We followed the original approach of LLaVA 1.5 (Liu et al., 2023b), we used the complete pre-training dataset of LLaVA 1.5 during the first stage of pre-training. We also keeping the same hyperparameter settings with LLaVA 1.5 . Our experiments were conducted using 16 NVIDIA A100 GPUs with 80G of memory. We used Deepspeed (Rajbhandari et al., 2020) for Hal-Evaluator and Hal-VL, with a batch size of 64 on a single GPU.
C.2. Setting for GPT-4 Annotation of AFHA
@@
C.3. Setting for GPT-4 Filtering of AFHA
@@
C.4. Experiment Setting of Chain-of-Thought for Discriminative Evaluation.
@@
Table 9. Evaluation results for different MLLMs on MMHal-Bench.
Figure 8. Evaluation results on Hal-Eval-Dis after fine-tuning Hal-Evaluator using SFT data of varying sizes. We have calculated the average F1 value across all types of hallucinations.
Method Overall Hallucination Score in Each Question Type $\uparrow$
@@
[…]
Table 10. Evaluation results for different MLLMs on MMHal-Bench.
@@
Table 11. Generative Hallucination Evaluation for Hal-VL and LLaVA 1.5.
Method MME MMBench MM-Vet SEED-Bench
BLIP-2 (Li et al., 2023c) 1293.84 - 22.4 46.4
mPLUG-Owl (Ye et al., 2023) 967.34 46.6 - 34.0
InstructBLIP (Dai et al., 2023) 1212.82 36.0 26.2 53.4
Otter (Li et al., 2023e) 1292.26 48.3 24.6 32.9
Qwen-VL-Chat (Bai et al., 2023b) 1487.58 60.6 - 58.2
LLaVA (Liu et al., 2023e) 502.82 36.2 28.1 33.5
MiniGPT-4 (Zhu et al., 2023a) 581.67 23.0 22.1 42.8
LLaVA-1.5 (Liu et al., 2023c) 1510.70 64.3 30.5 58.6
Hal-VL 1620.10 66.3 31.3 59.4
Table 12. Zero-shot multi-modal evaluation on multi-modal benchmarks including MME (Fu et al., 2023), MMBench (Liu et al., 2023a), MM-Vet (Yu et al., 2023), SEED-Bench (Li et al., 2023d). The overall scores are reported for evaluation. For MMBench, we report test results.
Model In-domain Out-of-domain Length
LLaVA1.5 23.7 58.8 10.6 7.0 55.7 30.0 48.4 11.6 10.2 49.5 10.3
LLaVA1.5 42.2 13.0 3.6 41.4 44.6 34.6 8.8 2.7 54.3 46.4 84.5
@@
Hal-VL 31.3 22.9 23.8 22.2 64.1 27.8 17.7 22.2 32.5 56.9 85.5
Table 13. Generative Hallucination Evaluation for Hal-VL and LLaVA 1.5.
Appendix D More Experiments
D.1. More result of Hal-VL.
In this section, we provide detailed evaluations of Hal-VL on more general benchmarks and Hal-Eval (Table LABEL:tab:dis_halvl and Table 13). As indicated in the Table 6, we found that Hal-VL achieved significant advantages on the vast majority of these benchmarks. This demonstrates its robustness and generalizability across different tasks and datasets, underlining its effectiveness in handling hallucination problems in LVLMs.
@@ Appendix D More Experiments
POPE (Li et al., 2023a) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
NOPE (Lovenia et al., 2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
CIEM (Hu et al., 2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
M-HalDetect (Gunjal et al., 2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
GAVIE (Liu et al., 2023f) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
FAITHScore (Jing et al., 2023) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
HaELM (Wang et al., 2023c) $\times$ $\times$ $\times$ $\times$ - - - - $\times$
MMHal-Bench (Sun et al., 2023a) $\times$ $\times$ $\times$ $\times$ - - - - $\times$
AMBER (Wang et al., 2023a) $\times$ $\times$ $\times$ $\times$
@@
Table 14. Comparison of Hallucination Evaluation Benchmarks for LVLMs.
D.2. Analysis of Different Types of Hallucinations Based on GPT-4
We further investigated the proportion of different types of hallucinations present within the output of LVLMs. As illustrated in Figure 2, we collected 5,000 image-caption pairs from COCO [26]. These were described by mPLUG-owl [50] and LLaVA [30], respectively. Subsequently, we provided both the ground truth image descriptions and the model-generated descriptions to GPT-4 [35]. We prompted it to evaluate whether these descriptions included hallucinations and to categorize them based on Object, Attribute, Relationship, and Event hallucinations.
Using the definitions of different types of hallucinations given in Section 2, we asked GPT-4 to first determine whether the model’s output contained hallucinations. If hallucinations were present, GPT-4 was instructed to classify them according to the definitions of different types of hallucinations. We tabulated the proportions of different types of hallucinations at varying description lengths. We found a significant increase in the proportion of event hallucinations as the length of the description extended.
D.3. Impact of SFT Data Scales for Hal-Evaluator
D.4. Effectiveness of Hall-evaluator for Mitigating Generative Hallucination.
The evaluation model Hal-Evaluator, utilized for generative hallucination assessment, can not only detect and evaluate model hallucinations but also modify the hallucinatory content within model outputs to aid in hallucination elimination. To validate the effectiveness of Hal-Evaluator in eliminating hallucinations, we conducted the following experiment: we used the output from MiniGPT-4 and LLaVA as the input for Hal-Evaluator, allowing Hal-Evaluator to detect and rectify any hallucinatory content. Subsequently, we re-evaluated the corrected output with MMHal-Bench (Sun et al., 2023a) as shown in Table 9. Our findings indicate that Hal-Evaluator can effectively eliminate model hallucinations.
D.5. Effectiveness of Hall-evaluator for Mitigating Generative Hallucination.
The evaluation model Hal-Evaluator, utilized for generative hallucination assessment, can not only detect and evaluate model hallucinations but also modify the hallucinatory content within model outputs to aid in hallucination elimination. To validate the effectiveness of Hal-Evaluator in eliminating hallucinations, we conducted the following experiment: we used the output from MiniGPT-4 and LLaVA as the input for Hal-Evaluator, allowing Hal-Evaluator to detect and rectify any hallucinatory content. Subsequently, we re-evaluated the corrected output with MMHal-Bench (Sun et al., 2023a) as shown in Table 9. Our findings indicate that Hal-Evaluator can effectively eliminate model hallucinations.
D.6. Compared with Other Hallucination Benchmarks
@@
Figure 9. The image shows the prompt we used to elicit hallucination annotations from GPT-4.
Figure 10. The figure illustrates the tailored prompt designed to guide GPT-4 in distinguishing and discarding labels that are inconsistent with the predefined criteria of hallucination types or are marked inaccurately.
Figure 11. The provided image displays an example utilized in the construction of the instruction tuning dataset, which we refer to as Hal-Data-SFT. The highlighted portions in green represent the elements that the model needs to predict and compute loss during training.
D.7. Demo for Hal-Evaluator